What number would you add to the equation below to complete the square?
x2 - 11x = 0

Answer by mananth(16946)

x^2-11x +c =0
a^2-2*a*b+b^2=0
x^2 -2 *1x*11/2 +c=0
C has to be the square off 11/2
= 121 /4
Ananth